OpenStudy (anonymous):

@StudyGurl14- During his first year of life, Avery gained 12 3/4 lb. During his second year of life, Avery gained 40% less weight than he did during his first year of life. How much weight did Avery gain during his second year of life? Express your answer as a mixed number in simplest form.

OpenStudy (studygurl14):

\(\large 12\frac{3}{4}-40\%(12\frac{3}{4})\)= the answer

OpenStudy (studygurl14):

\(\large 12\frac{3}{4}\rightarrow\frac{51}{4}\) \(40\%\rightarrow\large\frac{40}{100}\rightarrow\frac{2}{5}\) \(\large\frac{51}{4}-\frac{2}{5}(\frac{51}{4})\)

OpenStudy (studygurl14):

\(\Large\frac{51}{4}-\frac{102}{20}\) \(\Large\frac{51}{4}(\frac{5}{5})-\frac{102}{20}\rightarrow\frac{255}{20}-\frac{102}{20}=\tt the~answer\)

OpenStudy (studygurl14):

\(\large 255 - 102=153\) \(\Large\frac{153}{20}\), which simplifies into \(\Large 7\frac{13}{20}\)

OpenStudy (anonymous):

thnx